# Crystallization of discrete $N$-particle systems at high temperature

### v1

@@ 4.1. Preliminary lemmas
satisfies the LLN, in the sense of Definition˜3.2. In particular, if we let $m_{1},m_{2},\cdots$ be the moments of $\mu$, then
@@
Since $\{\mathbb{P}_{N}\}_{N\geq 1}$ satisfies the LLN, Theorem˜3.4 implies HT-appropriateness. In particular, we have the following convergence of Taylor coefficients of the logarithms:
@@
and (43) is valid whenever $\Re z\gg 0$. Next, we want to consider the asymptotic expansion of both sides in the region $|\arg z|<\pi/2$, as $z\to\infty$, in the sense of Eq.˜32. The LHS is the same because it is an analytic function, while the asymptotic expansion of the RHS can be deduced from Lemmas˜4.4 and 4.5. We end up with the following equality, which should be interpreted as an equality between formal power series in $z^{-1}$:
@@
It is easy to show that the empirical measures (36) corresponding to the sequence $\big\{\lambda_{n}^{(N)}\big\}_{N\geq 1}$ satisfy the LLN in the sense of Definition˜3.2, with corresponding $m_{1},m_{2},\cdots\in\mathbb{R}$ given by
@@ 5. Characteristic functions of the alpha, beta and Plancherel Jacobi operators
Let us specialize to three multiparameter families of Jacobi matrices, that will be labeled by “alpha”, “beta” and “planch”, in connection to ˜2.3. The Jacobi matrices $J^{(\text{alpha})}$ and $J^{(\text{planch})}$ will be infinite, while $J^{(\text{beta})}$ will be finite.
@@
Let $\gamma,\eta>0$, $c\in(0,1)$. The entire functions $F^{\text{alpha}}(z)$ and $F^{\text{planch}}(z)$, depending on the previous parameters and defined in Theorem 5.4, are of finite order.
@@ 6. The inverse of the formal Laplace transform
Immediately follows from Eq.˜63. ∎
@@
By the assumption of $G(z)$, we have $\lim_{z\to+\infty}{z\cdot\big(\ln{G(z)}\big)^{\prime}}=0$. We claim that the same is true of $z$ times the summation in the RHS of Eq.˜67. Indeed, if $z>0$ and $k$ is large enough so that $\ell_{k+1}<0$ and $|\ell_{k+1}+k|<\text{const}_{1}/k$, for some positive constant, then
@@
Since $\ell_{k+1}=-k+O\left(k^{-N}\right)$, for all $N\geq 1$, then the sums $\sum_{k=0}^{\infty}\Big(\big(\ell_{k+1}\big)^{n}-\big(-k\big)^{n}\Big)$ converge, for all $n\geq 1$. By expanding the RHS of Eq.˜68 around $z=\infty$, we obtain
@@
Finally, by Lemma˜6.1 and Eq.˜69, we deduce that
@@
Let $\gamma,\eta>0$ be arbitrary. Let $J^{(\text{planch})}$ be the Jacobi matrix that depends on these parameters, defined by Eq.˜53, and let $\ell_{1}^{(\text{planch})}>\ell_{2}^{(\text{planch})}>\dots$ be its eigenvalues. Then
@@
We will apply Theorem˜6.2 to $G(z)={}_{1}F_{1}\left(\gamma;z;-\eta\right)$. We need to check the conditions of that theorem. The fact that $F(z)=\frac{1}{\Gamma(z)}\cdot G(z)$ is entire and of finite order is proved in Lemma˜5.5, while the fact that the zeroes of $F(z)$ are simple and satisfy the desired asymptotic relation follows from Lemmas˜5.1 and 5.4. Finally, it is easy to verify from the power series definition of $_{1}F_{1}(\gamma;z;-\eta)$ that this function has a complete asymptotic expansion, as $z\to\infty$, beginning with $_{1}F_{1}(\gamma;z;-\eta)=1+O(z^{-1})$, as needed. ∎
@@
The statement of the theorem is quite similar to that of Theorem˜6.2. In fact, the only difference in the assumptions of the theorem is that now we have the asymptotic relation $\ell_{k}=-k+O\big(k^{-N}\big)$ instead of $\ell_{k+1}=-k+O\big(k^{-N}\big)$. The proof of the present theorem is similar; in fact, we will only point out the places in the proof where we need modifications.
As in the proof of Theorem˜6.2, we find
@@
where $R(z):=Q^{\prime}(z)-\gamma+\sum_{k=1}^{\infty}{\big\{\frac{1}{\ell_{k}}+\frac{1}{k}\big\}}$. Using the asymptotic relation $\ell_{k}=-k+O\left(k^{-N}\right)$, for any $N\geq 1$, as $k\to\infty$, as in the proof of Theorem˜6.2, we verify $\lim_{z\to\infty}{(m-1+zR(z))}=0$. This limit implies that the polynomial m1+zR(z)m-1+zR(z) is identically zero, therefore
@@
The proof then proceeds along the same lines as the proof of Theorem˜6.2. ∎
@@
Let $\gamma,\eta>0$, $c\in(0,1)$ be arbitrary. Let $J^{(\text{alpha})}$ be the Jacobi matrix that depends on these parameters, defined by Eq.˜51, and let $\ell_{1}^{(\text{alpha})}>\ell_{2}^{(\text{alpha})}>\dots$ be the eigenvalues of $J^{(\text{alpha})}$. If $\ell_{k}^{(\text{alpha})}=-k+O(k^{-N})$, for any $N>0$, as $k\to\infty$, then
@@
The first equality is Pfaff’s transformation, namely Lemma˜5.2. The second equality results from Theorem˜6.4 applied to $G(z)=(1-c)^{\gamma}\cdot{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)$, so it only remains to verify the assumptions. The function $F(z)=\frac{1}{\Gamma(z)}\cdot G(z)$ is entire and of finite order, due to Lemma˜5.5. The zeroes of $F(z)$ are simple and satisfy the desired asymptotic relations because of Lemma˜5.1, Theorem˜5.4, and the assumption of the corollary. Finally, $_{2}F_{1}\left(\gamma,z-\eta;z;c\right)$ is known to have an asymptotic expansion, as $z\to\infty$, that begins with $_{2}F_{1}\left(\gamma,z-\eta;z;c\right)\sim(1-c)^{-\gamma}\cdot\big(1+O(z^{-1})\big)$, see e.g. [Tem03], already cited in the proof of Theorem˜5.4. Therefore, $(1-c)^{\gamma}\cdot{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)\sim 1+O(z^{-1})$, as required. ∎
@@
Let $M\in\mathbb{Z}_{\geq 1}$, $c\in(0,1)$ be arbitrary. Let $J^{(\text{beta})}$ be the symmetric tridiagonal $M\times M$ matrix that depends on these parameters and is defined by Eq.˜52. The eigenvalues of $J^{(\text{beta})}$ or, equivalently, the roots of the monic degree $M$ polynomial $F^{\text{beta}}(z)$ are real; let us denote them $\ell_{1}^{(\text{beta})},\dots,\ell_{M}^{(\text{beta})}$. Then
@@ 7. Densities of the limits of pure Jack measures
In this section, as the main application of our main Theorem˜4.1, we find densities for the limits of empirical measures of pure Jack measures, i.e., for $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$, $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ and $\mu_{\operatorname{Planch}}^{\gamma;\eta}$, namely the measures from Theorem˜3.6, Theorem˜3.7 and Theorem˜3.8, respectively.
@@
Let $\gamma,\eta>0$ be arbitrary. Let $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ be the probability measure from Theorem˜3.8. Let $\ell_{1}^{(\text{planch})}>\ell_{2}^{(\text{planch})}>\cdots$ be the eigenvalues of $J^{(\text{planch})}$ or, equivalently, the zeroes of $F^{\text{planch}}(z)$ [^7] The parameters $\gamma,\eta$ that go into the definitions of the Jacobi operator $J^{(\text{planch})}$ and the meromorphic function $F^{\text{planch}}(z)$ are the same as the ones in $\mu_{\operatorname{Planch}}^{\gamma;\eta}$.; see Eqs.˜53 and 5.4 for the definitions. Then $\gamma\geq\ell_{1}^{(\text{planch})}$ and $\ell_{k}^{(\text{planch})}\geq 1+\ell_{k+1}^{(\text{planch})}$, for all $k\geq 1$. Moreover, $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ is absolutely continuous and has density
@@
By Theorem˜3.8, the measure $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ has finite moments of all orders and, consequently, all its quantized $\gamma$-cumulants are well-defined; moreover, they equal $\kappa_{n}=\eta\cdot\delta_{n,1}$. Then Eq.˜23 gives
@@
Denote by $M(s):=1+\sum_{n=1}^{\infty}{\int_{\mathbb{R}}{x^{n}\mu_{\operatorname{Planch}}^{\gamma;\eta}(\operatorname{d\!}x)}\cdot s^{n}}$ the formal moment generating function of $\mu_{\operatorname{Planch}}^{\gamma;\eta}$. By Theorem˜4.1 and Corollary˜6.3, we deduce the following equality of formal power series:
@@
Then $1-\ell_{k}^{(\text{planch})}\leq-\ell_{k+1}^{(\text{planch})}$, for all $k\geq 0$, so that all intervals $\big[1-\ell_{k}^{(\text{planch})},\,-\ell_{k+1}^{(\text{planch})}\big]$, for $k\geq 0$, are well-defined and disjoint (if $1-\ell_{k}^{(\text{planch})}=-\ell_{k+1}^{(\text{planch})}$, for some $k$, then we can remove the corresponding indicator function from the density). Moreover, the function defined by Eq.˜71 is the density of a probability measure, since it integrates to:
@@
where for the fourth equality, we used $\lim_{K\to\infty}{\big(\ell_{K+1}^{(\text{planch})}+K\big)}=0$, as given by Lemma˜5.1. Next, note that $\int_{\mathbb{R}}{e^{itx}f_{\operatorname{Planch}}^{\gamma;\eta}(x)dx}$ is equal to the RHS of Eq.˜72. Since the characteristic function uniquely determines the underlying probability measure, it would follow that $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ is absolutely continuous with density $f_{\operatorname{Planch}}^{\gamma;\eta}(x)$, as desired.
It remains to prove Eq.˜73. If we set
@@
then $\int_{\mathbb{R}}{e^{itx}g_{\operatorname{Planch}}^{\gamma;\eta}(x)\operatorname{d\!}x}$ equals the RHS of Eq.˜72. This means that $g_{\operatorname{Planch}}^{\gamma;\eta}(x)$ is a density of the probability measure $\mu_{\operatorname{Planch}}^{\gamma;\eta}$. For the sake of contradiction, assume $\ell_{1}^{(\text{planch})}<-\gamma$. Since $\ell_{1}^{(\text{planch})}<1-\ell_{k}^{(\text{planch})}$ and $\ell_{1}^{(\text{planch})}<-\ell_{k+1}^{(\text{planch})}$, for all $k\geq 1$, it is easy to check that $g_{\operatorname{Planch}}^{\gamma;\eta}\big(-\ell_{1}^{(\text{planch})}+\epsilon\big)<0$, for sufficiently small $\epsilon>0$. But this is impossible, therefore $\ell_{1}^{(\text{planch})}\geq-\gamma$, which proves the first equality in (73). Similarly, $\ell_{k+1}^{(\text{planch})}\geq 1-\ell_{k}^{(\text{planch})}$, for all $k\geq 1$, by induction on $k$. This ends the proof of (73) and the theorem. ∎
@@
Let $\gamma,\eta>0$ and $c\in(0,1)$ be arbitrary. Let $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ be the probability measure from Theorem˜3.6. Let $\ell_{1}^{(\text{alpha})}>\ell_{2}^{(\text{alpha})}>\cdots$ be the eigenvalues of $J^{(\text{alpha})}$ or, equivalently, the zeroes of $F^{\text{alpha}}(z)$. Moreover, assume that $\ell_{k}^{(\text{alpha})}=-k+O(k^{-N})$, for any $N\geq 1$, as $k\to\infty$. [^8] This assumption holds when $\gamma=\eta$ and we believe that it holds for all $\gamma,\eta>0$. Then $1\geq\ell_{1}^{(\text{alpha})}$ and $\ell_{k}^{(\text{alpha})}\geq 1+\ell_{k+1}^{(\text{alpha})}$, for all $k\geq 1$. Moreover, $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ is absolutely continuous and has density
@@
By Theorem˜3.8, the quantized $\gamma$-cumulants of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ are $\kappa_{n}=\eta\left(\frac{c}{1-c}\right)^{n}$, thus Eq.˜23 gives
@@
where we have used Lemma˜5.2. The advantage of the last two formulas is that the corresponding series define analytic functions, for any $c\in(0,1)$. Next, if we let $M(s)$ be the formal moment generating function of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$, then Theorems˜4.1 and 6.5 imply the following equality of formal power series:
@@
The remaining of the proof then follows the proof of Theorem˜7.1, starting by computing that the characteristic function of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ is
@@
Let $M\in\mathbb{Z}_{\geq 1}$, $c\in(0,1)$. Let $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ be the probability measure from Theorem˜3.7. Let $\ell_{1}^{(\text{beta})}\geq\dots\geq\ell_{M}^{(\text{beta})}$ be the eigenvalues of $J^{(\text{beta})}$ or, equivalently, the roots of $F^{\text{beta}}(z)$. Then $\gamma\geq\ell_{1}^{(\text{beta})}$ and $\ell_{k}^{(\text{beta})}\geq 1+\ell_{k+1}^{(\text{beta})}$, for all $1\leq k\leq M-1$; in particular, the eigenvalues are pairwise distinct. Moreover, $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ has density
@@
By Theorem˜3.7, the quantized $\gamma$-cumulants of $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ are $\kappa_{n}=(-1)^{n-1}Mc^{n}$. Then Eq.˜23 gives
@@
Then, if we let $\mathfrak{M}(s)$ be the formal moment generating function of $\mu_{\operatorname{Beta}}^{\gamma;c;M}$, and due to the already proved Eq.˜70, we deduce from Theorem˜4.1 the following equality of power series:
@@
From here, the proof mimics the proof of Theorem˜7.1. In fact, the previous equality gives

### v2

@@ 4.2. Proof of Theorem 4
satisfies the LLN, in the sense of Section˜3.1. In particular, if we let $m_{1},m_{2},\cdots$ be the moments of $\mu$, then
@@
Since $\{\mathbb{P}_{N}\}_{N\geq 1}$ satisfies the LLN, Section˜3.1 implies HT-appropriateness. In particular, we have the following convergence of Taylor coefficients of the logarithms:
@@
and (43) is valid whenever $\Re z\gg 0$. Next, we want to consider the asymptotic expansion of both sides in the region $|\arg z|<\pi/2$, as $z\to\infty$, in the sense of Equation˜32. The LHS is the same because it is an analytic function, while the asymptotic expansion of the RHS can be deduced from Lemmas˜4.4 and 4.5. We end up with the following equality, which should be interpreted as an equality between formal power series in $z^{-1}$:
@@
It is easy to show that the empirical measures (36) corresponding to the sequence $\big\{\lambda_{n}^{(N)}\big\}_{N\geq 1}$ satisfy the LLN in the sense of Section˜3.1, with corresponding $m_{1},m_{2},\cdots\in\mathbb{R}$ given by
@@ 5. Characteristic functions of the alpha, beta and Plancherel Jacobi operators
Let us specialize to three multiparameter families of Jacobi matrices, that will be labeled by “alpha”, “beta” and “planch”, in connection to Section˜2.5. The Jacobi matrices $J^{(\text{alpha})}$ and $J^{(\text{planch})}$ will be infinite, while $J^{(\text{beta})}$ will be finite.
@@
Let $\gamma,\eta>0$, $c\in(0,1)$. The entire functions $F^{\text{alpha}}(z)$ and $F^{\text{planch}}(z)$, depending on the previous parameters and defined in Theorem 5, are of finite order.
@@ 6. The inverse of the formal Laplace transform
Immediately follows from Equation˜63. ∎
@@
By the assumption of $G(z)$, we have $\lim_{z\to+\infty}{z\cdot\big(\ln{G(z)}\big)^{\prime}}=0$. We claim that the same is true of $z$ times the summation in the RHS of Equation˜67. Indeed, if $z>0$ and $k$ is large enough so that $\ell_{k+1}<0$ and $|\ell_{k+1}+k|<\text{const}_{1}/k$, for some positive constant, then
@@
Since $\ell_{k+1}=-k+O\left(k^{-N}\right)$, for all $N\geq 1$, then the sums $\sum_{k=0}^{\infty}\Big(\big(\ell_{k+1}\big)^{n}-\big(-k\big)^{n}\Big)$ converge, for all $n\geq 1$. By expanding the RHS of Equation˜68 around $z=\infty$, we obtain
@@
Finally, by Lemma˜6.1 and Equation˜69, we deduce that
@@
Let $\gamma,\eta>0$ be arbitrary. Let $J^{(\text{planch})}$ be the Jacobi matrix that depends on these parameters, defined by Equation˜53, and let $\ell_{1}^{(\text{planch})}>\ell_{2}^{(\text{planch})}>\dots$ be its eigenvalues. Then
@@
We will apply Section˜6 to $G(z)={}_{1}F_{1}\left(\gamma;z;-\eta\right)$. We need to check the conditions of that theorem. The fact that $F(z)=\frac{1}{\Gamma(z)}\cdot G(z)$ is entire and of finite order is proved in Lemma˜5.5, while the fact that the zeroes of $F(z)$ are simple and satisfy the desired asymptotic relation follows from Lemmas˜5.1 and 5. Finally, it is easy to verify from the power series definition of $_{1}F_{1}(\gamma;z;-\eta)$ that this function has a complete asymptotic expansion, as $z\to\infty$, beginning with $_{1}F_{1}(\gamma;z;-\eta)=1+O(z^{-1})$, as needed. ∎
@@
The statement of the theorem is quite similar to that of Section˜6. In fact, the only difference in the assumptions of the theorem is that now we have the asymptotic relation $\ell_{k}=-k+O\big(k^{-N}\big)$ instead of $\ell_{k+1}=-k+O\big(k^{-N}\big)$. The proof of the present theorem is similar; in fact, we will only point out the places in the proof where we need modifications.
As in the proof of Section˜6, we find
@@
where $R(z):=Q^{\prime}(z)-\gamma+\sum_{k=1}^{\infty}{\big\{\frac{1}{\ell_{k}}+\frac{1}{k}\big\}}$. Using the asymptotic relation $\ell_{k}=-k+O\left(k^{-N}\right)$, for any $N\geq 1$, as $k\to\infty$, as in the proof of Section˜6, we verify $\lim_{z\to\infty}{(m-1+zR(z))}=0$. This limit implies that the polynomial m1+zR(z)m-1+zR(z) is identically zero, therefore
@@
The proof then proceeds along the same lines as the proof of Section˜6. ∎
@@
Let $\gamma,\eta>0$, $c\in(0,1)$ be arbitrary. Let $J^{(\text{alpha})}$ be the Jacobi matrix that depends on these parameters, defined by Equation˜51, and let $\ell_{1}^{(\text{alpha})}>\ell_{2}^{(\text{alpha})}>\dots$ be the eigenvalues of $J^{(\text{alpha})}$. If $\ell_{k}^{(\text{alpha})}=-k+O(k^{-N})$, for any $N>0$, as $k\to\infty$, then
@@
The first equality is Pfaff’s transformation, namely Lemma˜5.2. The second equality results from Section˜6 applied to $G(z)=(1-c)^{\gamma}\cdot{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)$, so it only remains to verify the assumptions. The function $F(z)=\frac{1}{\Gamma(z)}\cdot G(z)$ is entire and of finite order, due to Lemma˜5.5. The zeroes of $F(z)$ are simple and satisfy the desired asymptotic relations because of Lemma˜5.1, Section˜5, and the assumption of the corollary. Finally, $_{2}F_{1}\left(\gamma,z-\eta;z;c\right)$ is known to have an asymptotic expansion, as $z\to\infty$, that begins with $_{2}F_{1}\left(\gamma,z-\eta;z;c\right)\sim(1-c)^{-\gamma}\cdot\big(1+O(z^{-1})\big)$, see e.g. [Tem03], already cited in the proof of Section˜5. Therefore, $(1-c)^{\gamma}\cdot{}_{2}F_{1}\left(\gamma,z-\eta;z;c\right)\sim 1+O(z^{-1})$, as required. ∎
@@
Let $M\in\mathbb{Z}_{\geq 1}$, $c\in(0,1)$ be arbitrary. Let $J^{(\text{beta})}$ be the symmetric tridiagonal $M\times M$ matrix that depends on these parameters and is defined by Equation˜52. The eigenvalues of $J^{(\text{beta})}$ or, equivalently, the roots of the monic degree $M$ polynomial $F^{\text{beta}}(z)$ are real; let us denote them $\ell_{1}^{(\text{beta})},\dots,\ell_{M}^{(\text{beta})}$. Then
@@ 7. Densities of the limits of pure Jack measures
In this section, as the main application of our main Section˜4, we find densities for the limits of empirical measures of pure Jack measures, i.e., for $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$, $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ and $\mu_{\operatorname{Planch}}^{\gamma;\eta}$, namely the measures from Section˜3.2.1, Section˜3.2.2 and Section˜3.2.3, respectively.
@@
Let $\gamma,\eta>0$ be arbitrary. Let $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ be the probability measure from Section˜3.2.3. Let $\ell_{1}^{(\text{planch})}>\ell_{2}^{(\text{planch})}>\cdots$ be the eigenvalues of $J^{(\text{planch})}$ or, equivalently, the zeroes of $F^{\text{planch}}(z)$ [^7] The parameters $\gamma,\eta$ that go into the definitions of the Jacobi operator $J^{(\text{planch})}$ and the meromorphic function $F^{\text{planch}}(z)$ are the same as the ones in $\mu_{\operatorname{Planch}}^{\gamma;\eta}$.; see Equations˜53 and 5 for the definitions. Then $\gamma\geq\ell_{1}^{(\text{planch})}$ and $\ell_{k}^{(\text{planch})}\geq 1+\ell_{k+1}^{(\text{planch})}$, for all $k\geq 1$. Moreover, $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ is absolutely continuous and has density
@@
By Section˜3.2.3, the measure $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ has finite moments of all orders and, consequently, all its quantized $\gamma$-cumulants are well-defined; moreover, they equal $\kappa_{n}=\eta\cdot\delta_{n,1}$. Then Equation˜23 gives
@@
Denote by $M(s):=1+\sum_{n=1}^{\infty}{\int_{\mathbb{R}}{x^{n}\mu_{\operatorname{Planch}}^{\gamma;\eta}(\operatorname{d\!}x)}\cdot s^{n}}$ the formal moment generating function of $\mu_{\operatorname{Planch}}^{\gamma;\eta}$. By Section˜4 and Section˜6, we deduce the following equality of formal power series:
@@
Then $1-\ell_{k}^{(\text{planch})}\leq-\ell_{k+1}^{(\text{planch})}$, for all $k\geq 0$, so that all intervals $\big[1-\ell_{k}^{(\text{planch})},\,-\ell_{k+1}^{(\text{planch})}\big]$, for $k\geq 0$, are well-defined and disjoint (if $1-\ell_{k}^{(\text{planch})}=-\ell_{k+1}^{(\text{planch})}$, for some $k$, then we can remove the corresponding indicator function from the density). Moreover, the function defined by Equation˜71 is the density of a probability measure, since it integrates to:
@@
where for the fourth equality, we used $\lim_{K\to\infty}{\big(\ell_{K+1}^{(\text{planch})}+K\big)}=0$, as given by Lemma˜5.1. Next, note that $\int_{\mathbb{R}}{e^{itx}f_{\operatorname{Planch}}^{\gamma;\eta}(x)dx}$ is equal to the RHS of Equation˜72. Since the characteristic function uniquely determines the underlying probability measure, it would follow that $\mu_{\operatorname{Planch}}^{\gamma;\eta}$ is absolutely continuous with density $f_{\operatorname{Planch}}^{\gamma;\eta}(x)$, as desired.
It remains to prove Equation˜73. If we set
@@
then $\int_{\mathbb{R}}{e^{itx}g_{\operatorname{Planch}}^{\gamma;\eta}(x)\operatorname{d\!}x}$ equals the RHS of Equation˜72. This means that $g_{\operatorname{Planch}}^{\gamma;\eta}(x)$ is a density of the probability measure $\mu_{\operatorname{Planch}}^{\gamma;\eta}$. For the sake of contradiction, assume $\ell_{1}^{(\text{planch})}<-\gamma$. Since $\ell_{1}^{(\text{planch})}<1-\ell_{k}^{(\text{planch})}$ and $\ell_{1}^{(\text{planch})}<-\ell_{k+1}^{(\text{planch})}$, for all $k\geq 1$, it is easy to check that $g_{\operatorname{Planch}}^{\gamma;\eta}\big(-\ell_{1}^{(\text{planch})}+\epsilon\big)<0$, for sufficiently small $\epsilon>0$. But this is impossible, therefore $\ell_{1}^{(\text{planch})}\geq-\gamma$, which proves the first equality in (73). Similarly, $\ell_{k+1}^{(\text{planch})}\geq 1-\ell_{k}^{(\text{planch})}$, for all $k\geq 1$, by induction on $k$. This ends the proof of (73) and the theorem. ∎
@@
Let $\gamma,\eta>0$ and $c\in(0,1)$ be arbitrary. Let $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ be the probability measure from Section˜3.2.1. Let $\ell_{1}^{(\text{alpha})}>\ell_{2}^{(\text{alpha})}>\cdots$ be the eigenvalues of $J^{(\text{alpha})}$ or, equivalently, the zeroes of $F^{\text{alpha}}(z)$. Moreover, assume that $\ell_{k}^{(\text{alpha})}=-k+O(k^{-N})$, for any $N\geq 1$, as $k\to\infty$. [^8] This assumption holds when $\gamma=\eta$ and we believe that it holds for all $\gamma,\eta>0$. Then $1\geq\ell_{1}^{(\text{alpha})}$ and $\ell_{k}^{(\text{alpha})}\geq 1+\ell_{k+1}^{(\text{alpha})}$, for all $k\geq 1$. Moreover, $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ is absolutely continuous and has density
@@
By Section˜3.2.3, the quantized $\gamma$-cumulants of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ are $\kappa_{n}=\eta\left(\frac{c}{1-c}\right)^{n}$, thus Equation˜23 gives
@@
where we have used Lemma˜5.2. The advantage of the last two formulas is that the corresponding series define analytic functions, for any $c\in(0,1)$. Next, if we let $M(s)$ be the formal moment generating function of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$, then Sections˜4 and 6 imply the following equality of formal power series:
@@
The remaining of the proof then follows the proof of Section˜7, starting by computing that the characteristic function of $\mu_{\operatorname{Alpha}}^{\gamma;c;\eta}$ is
@@
Let $M\in\mathbb{Z}_{\geq 1}$, $c\in(0,1)$. Let $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ be the probability measure from Section˜3.2.2. Let $\ell_{1}^{(\text{beta})}\geq\dots\geq\ell_{M}^{(\text{beta})}$ be the eigenvalues of $J^{(\text{beta})}$ or, equivalently, the roots of $F^{\text{beta}}(z)$. Then $\gamma\geq\ell_{1}^{(\text{beta})}$ and $\ell_{k}^{(\text{beta})}\geq 1+\ell_{k+1}^{(\text{beta})}$, for all $1\leq k\leq M-1$; in particular, the eigenvalues are pairwise distinct. Moreover, $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ has density
@@
By Section˜3.2.2, the quantized $\gamma$-cumulants of $\mu_{\operatorname{Beta}}^{\gamma;c;M}$ are $\kappa_{n}=(-1)^{n-1}Mc^{n}$. Then Equation˜23 gives
@@
Then, if we let $\mathfrak{M}(s)$ be the formal moment generating function of $\mu_{\operatorname{Beta}}^{\gamma;c;M}$, and due to the already proved Equation˜70, we deduce from Section˜4 the following equality of power series:
@@
From here, the proof mimics the proof of Section˜7. In fact, the previous equality gives